OpenStudy (anonymous):

divide the complex number :

OpenStudy (anonymous):

\[\frac{ 4-2i }{ 3i}\]

OpenStudy (anonymous):

multiply top and bottom either by \(i\) or \(-i\) your choice

OpenStudy (anonymous):

how exactly is that done?

OpenStudy (anonymous):

if you multiply by \(i\) you get as a first step \[\frac{4i-2i^2}{3i^2}\] as a first steo, then, since \(i^2=-1\) it is \[\frac{4i+2}{-3}\] as step 2 usually you skip the first step and go right to the second

OpenStudy (anonymous):

then to put the complex number in standard from of \(a+bi\) you write it as \[-\frac{2}{3}-\frac{4}{3}i\]

OpenStudy (anonymous):

they have common denominators

OpenStudy (anonymous):

yes, of course that is because it started as \[\frac{2+4i}{-3}\] but standard from is \(a+bi\) so you break it apart

OpenStudy (anonymous):

so the next step is to simplify correct